OpenStudy (anonymous):

ILL GIVE A MEDAL TO THE BEST ANSWER Mr.Hall is planting a circular flower garden made up of Texas Bluebonnets. It is 144π ft^2 in area. What is the radius of the garden? @jdoe0001 @whpalmer4

OpenStudy (hoblos):

\[area = \pi r^2 = 144 \pi\] so \[r^2 = \frac{ 144 \pi }{ \pi } = 144\] so what's r ?

OpenStudy (anonymous):

r=12 @hoblos

OpenStudy (hoblos):

correct